# Data-driven geometric parameter optimization for PD-GMRES

### v1

@@ 4.4 Classification and optimization of problematic matrices
Figure 4.6: Comparing performance of GMRES(100) and the previously optimized PD-GMRES against the specialized PD-GMRES found by optimizing the problematic training matrices. Normalizing around the optimized PD-GMRES the geometric means are: $0.8147$ for GMRES(100), $1$ for optimized PD-GMRES, $0.6676$ for the specialized PD-GMRES.
@@ 4.5 Evaluating a different parameter quadrant
All previous analysis focused on the $(\alpha_{p}<0,\alpha_{d}>0)$-quadrant. We now turn to the $(\alpha_{p}>0,\alpha_{d}>0)$-quadrant which is evaluated for $m_{\text{init}}=10$ and $m_{\text{init}}=30$ using the same initial choices for $m_{\min}$ and $m_{\text{step}}$ and the same quadtree size and resolution as before, but with $\alpha_{p}$ being evaluated in the interval $[0,40]$.
@@ 4.6 An algorithmic extension for PD-GMRES
All data presented here was calculated on an Intel i7-12700k CPU using MATLAB version R2019a.
@@ 5 Conclusions
We have presented a quadtree-based geometric method for solving two-dimensional optimization problems and have applied this method to the problem of finding optimal parameters for PD-GMRES, such that the runtime of PD-GMRES is minimized. Performing the optimization procedure on a set of training matrices we have found well-performing PD-GMRES parameters which were then validated on a set of test matrices. We have further confirmed the effectiveness of our optimization approach by applying it to a subset of training matrices which exhibited problematic behavior. This application has provided even better PD-GMRES performance and has shown that the quadtree method is a flexible tool for optimizing PD-GMRES parameters for sets of matrices. Employing the optimization procedure to a different parameter quadrant did not result in better performance. Attempts to find a simple classifier for the problematic matrices were unsuccessful. We have further introduced an extension of the PD-GMRES algorithm which uses an additional parameter to constrain the restart parameter in a fixed range. We have applied this modified algorithm to a specific matrix for which PD-GMRES has long runtime and have found improved performance.
We identify several directions for further investigations. As systems of linear equations are often preconditioned before solving, the effect of preconditioning on the evaluated matrix sets should be investigated. One specific question would be whether there exist any clear trends for how the PD-GMRES parameters should be changed when switching from an un-preconditioned system to the same system with a preconditioner. Further attempts could be made to classify the problematic matrices. As we have found well-performing PD-GMRES parameters for both the non-problematic and the problematic matrices, an efficient classifier would allow for an a-priori determination of which parameters should be used for a given matrix. However, it seems likely that there is no such universal classifier. Alternative criteria for the construction of the quadtree could be considered, as briefly outlined at the end of Section 3.1. The extended PD-GMRES algorithm using the additional parameter $m_{\max}$ should be further investigated. If some practical estimates for $m_{\max}$ can be found, the extended algorithm could be generally superior to the original. Here, matrices for which PD-GMRES performs very poorly should be of particular focus. It is also possible to apply the optimization procedure to PD-GMRES variants which make use of augmentation or deflation of the Krylov subspace to improve the rate of convergence [3].

### v2

@@ 4.4 Classification and optimization of problematic matrices
Figure 4.6: Comparing performance of GMRES(100) and the previously optimized PD-GMRES against the specialized PD-GMRES found by optimizing the problematic training matrices. Normalizing around the optimized PD-GMRES the geometric means are: $0.8147$ for GMRES(100), 11 for optimized PD-GMRES, $0.6676$ for the specialized PD-GMRES.
@@
4.5 Large matrix test
Practical problems often involve larger matrices than those evaluated so far and also require the use of preconditioners. We select a new set of matrices from [5], focusing on large matrices (roughly an order of magnitude larger than previous ones), which all relate to computational fluid dynamics problems. For all matrices in this set (cf. Table 4.8) GMRES($m$) does not converge within a reasonable time when not using a preconditioner, so preconditioning is necessary. We choose ILU-based preconditioners for all large matrices, using ILU(0) (zero-fill) for problems without zeros on the diagonal and ILUTP (thresholding and pivoting) [26] for the remaining ones. Here, we reduce the threshold parameter $\tau$ with increasing matrix size to keep the runtime of all matrices somewhat similar. By comparing our optimized PD-GMRES (cf. Section 4.2) to GMRES($m$) we find that, for a majority of the large matrices, GMRES($100$) performs better than PD-GMRES. This means that by our classification these matrices are considered problematic therefore suitable for applying our specialized PD-GMRES (cf. Section 4.4). Table 4.9 shows that the specialized PD-GMRES outperforms all considered GMRES($m$), even for non-problematic large matrices where the optimized PD-GMRES already performs well. The geometric mean of GMRES($100$) runtime ($m=100$ being the only value for which GMRES($m$) converges for all large matrices) relative to the runtime of the specialized PD-GMRES is $1.6006$.
Name n Non-zeros Condition Preconditioner
Goodwin_040 $17922$ $561677$ $2.56e+06$ ILUTP, $\tau=0.05$
Goodwin_054 $32510$ $1030878$ $6.30e+06$ ILUTP, $\tau=0.04$
Goodwin_071 $56021$ $1797934$ $1.43e+07$ ILUTP, $\tau=0.03$
Goodwin_095 $100037$ $3226066$ $3.43e+07$ ILUTP, $\tau=0.02$
Goodwin_127 $178437$ $5778545$ $8.20e+07$ ILUTP, $\tau=0.01$
venkat25 $62424$ $1717763$ $1.22e+08$ ILU(0)
venkat50 $62424$ $1717777$ $5.32e+08$ ILU(0)
raefsky3 $21200$ $1488768$ $5.36e+11$ ILU(0)
water_tank $60740$ $2035281$ $2.52e+09$ ILU(0)
poisson3Db $85623$ $2374949$ $1.66e+05$ ILU(0)
Table 4.8: Properties of the set of large matrices which are evaluated with preconditioners. All matrices are related to computational fluid dynamics problems. Since the Goodwin matrices have zeros on the diagonal, ILU(0) is not applicable and we choose ILUTP instead. We lower the ILUTP threshold $\tau$ with increasing matrix size to keep the PD-GMRES runtime balanced.
PD-GMRES GMRES($m$)
Name Optimized Specialized $m=10$ $m=20$ $m=30$ $m=50$ $m=100$
Goodwin_040 1.9403 0.3513 - - - 1.6612 0.6593
Goodwin_054 5.5042 0.7044 - - - - 0.9228
Goodwin_071 1.3334 0.4419 - - 3.5029 1.1785 0.7436
Goodwin_095 8.2934 1.4191 - - 10.294 5.1581 1.9296
Goodwin_127 46.832 3.8337 - - - - 8.0535
venkat25 0.5007 0.6679 0.7832 0.8158 0.8721 0.9988 1.1896
venkat50 0.7156 0.9495 1.1685 1.1585 1.2641 1.4927 1.8642
raefsky3 0.2240 0.1861 1.5778 0.4343 0.3713 0.3332 0.2732
water_tank 19.893 2.9135 - - - - 4.1406
poisson3Db 0.6422 0.6984 1.4529 1.1387 0.9429 0.9364 0.8939
Table 4.9: Runtimes for the set of large matrices using GMRES($m$) with $m=10,20,30,50,100$ and PD-GMRES with optimized and specialized parameters from previous sections. For each problem the best performing algorithm is marked in bold. Missing entries indicate failure to converge within $240$ seconds.
For a comparison of the total cost of solving the linear system, the cost of calculating the ILU-based preconditioners would have to be added. We did not include this cost in our comparison as the preconditioners were calculated once in advance and were also not specifically optimized beyond general suitability.
@@ 4.6 Evaluating a different parameter quadrant
All previous analysis focused on the $(\alpha_{p}<0,\alpha_{d}>0)$-quadrant. We now turn to the $(\alpha_{p}>0,\alpha_{d}>0)$-quadrant which is evaluated for $m_{\text{init}}=10$ and $m_{\text{init}}=30$ using the same training and test matrix sets, the same initial choices for $m_{\min}$ and $m_{\text{step}}$ and the same quadtree size and resolution as before, but with $\alpha_{p}$ being evaluated in the interval $[0,40]$.
@@ 4.7 An algorithmic extension for PD-GMRES
All data presented here was calculated using MATLAB version R2019a on an Intel i7-12700k CPU, except for Section 4.5 for which an AMD Ryzen 7 9800X3D CPU was used. This means that the runtime values shown in Table 4.9 cannot be directly compared to those in Table 4.3 or Table 4.6.
@@ 5 Conclusions
We have presented a quadtree-based geometric method for solving two-dimensional optimization problems and have applied this method to the problem of finding optimal parameters for PD-GMRES, such that the runtime of PD-GMRES is minimized. Performing the optimization procedure on a set of training matrices we have found well-performing PD-GMRES parameters which were then validated on a set of test matrices. We expect that these parameters will perform well for other matrices and could be considered candidates for standard parameters to be used for PD-GMRES. We have further confirmed the effectiveness of our optimization approach by applying it to a subset of training matrices which exhibited problematic behavior. This application has provided even better PD-GMRES performance and has shown that the quadtree method is a flexible tool for optimizing PD-GMRES parameters for sets of matrices. Moreover, a large matrix test has indicated the suitability of the obtained parameters for PD-GMRES in combination with preconditioning. Employing the optimization procedure to a different parameter quadrant did not result in better performance. Attempts to find a simple classifier for the problematic matrices were unsuccessful. We have further introduced an extension of the PD-GMRES algorithm which uses an additional parameter to constrain the restart parameter in a fixed range. We have applied this modified algorithm to a specific matrix for which PD-GMRES has long runtime and have found improved performance.
We identify several directions for further investigations. As systems of linear equations are often preconditioned before solving, the effect of preconditioning on the evaluated matrix sets should be further investigated. One specific question would be whether there exist any clear trends for how the PD-GMRES parameters should be changed when switching from an un-preconditioned system to the same system with a preconditioner. Further attempts could be made to classify the problematic matrices. As we have found well-performing PD-GMRES parameters for both the non-problematic and the problematic matrices, an efficient classifier would allow for an a-priori determination of which parameters should be used for a given matrix. However, it seems likely that there is no such universal classifier. Alternative criteria for the construction of the quadtree could be considered, as briefly outlined at the end of Section 3.1. The extended PD-GMRES algorithm using the additional parameter $m_{\max}$ should be further investigated. If some practical estimates for $m_{\max}$ can be found, the extended algorithm could be generally superior to the original. Here, matrices for which PD-GMRES performs very poorly should be of particular focus. It is also possible to apply the optimization procedure to PD-GMRES variants which make use of augmentation or deflation of the Krylov subspace to improve the rate of convergence [3].